Let x be a random variable that represents the commute time of teachers in a large urban area. If x is normally distributed with a mean of 46 minutes and a standard deviation of 5 minutes, find the commute time which is at the top of the bottom 25%.

Answer by Boreal(15235)

Need to find the z value for the 25th percentile.
It is -0.675
z=(x-mu)/sd
-0.675=(x-46)/5
-3.375=x-46
42.625 minutes=x
This is -0.675 sd below the mean, and that represents the 25th percentile.